OpenStudy (anonymous):

Solve the triangle. A = 48°, a = 32, b = 27 I am completely blanking on how to do this, so if you could please do a step by step process, that would be greatly appreciated :)

OpenStudy (shamil98):

solve for what?

OpenStudy (shamil98):

solve for c?

OpenStudy (anonymous):

Solve for B, C, and c.

OpenStudy (shamil98):

Is this a right triangle?

OpenStudy (anonymous):

It doesn't say, but I don't think it is; however, I'm not sure.

OpenStudy (shamil98):

is there a picture of it?

OpenStudy (anonymous):

No. Literally the only thing they give is exactly what I posted.

OpenStudy (shamil98):

triangles = 180 degrees 48 + B + C = 180 B + C = 132

OpenStudy (shamil98):

B = 132 - C C = 132 - B

OpenStudy (anonymous):

Use the cosine rule

OpenStudy (anonymous):

I am so sorry, but can you please refresh me on the cosine rule?

OpenStudy (anonymous):

http://img.docstoccdn.com/thumb/orig/122204151.png

OpenStudy (anonymous):

Just rearrange it for c^2

OpenStudy (anonymous):

Can you help me plug it in? I'm still confused. Also, isn't there then something with the arccos at the end?